Three metal rods of same lengths and same area of cross section having conductivities 1,2,3 units are connected in series.Then their effective conductivity will be

Grade:11

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

k1 = 1
k2 = 2
k3 = 3
in series
k (eq)_ = 1/k1+1/k2+1/k3
= 1/1+1/2+1/3 = 11/6
so, final answer will be 6/11
beacuse 1/ k (eq ) = 11/6

Regards
Khimraj
4 years ago
 
can u remember to find out the effective resistance in the case of parallel combination of resister . 
the same formula will going to be used in this case also . 
k1 = 1 
k2 = 2 
k3 = 3
in series 
k (eq)_ = 1/k1+1/k2+1/k3 
= 1/1+1/2+1/3 = 11/6 
so, final answer will be 6/11 
beacuse 1/ k (eq ) = 11/6 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..